CHARACTERIZATION OF THE X̃ ^{2}A_{1} (0,0,0) GROUND VIBRONIC STATE OF CH_{2}^{+} BY PFI-ZEKE PHO-TOELECTRON SPECTROSCOPY

\underline{S. WILLITSCH}, L. IMBACH, and F. MERKT, Physical Chemistry, ETH Zurich, CH-8093 Zurich.

The X̃ ^{2}A_{1} (0,0,0) ground vibronic state of CH_{2}^{+} has been studied by pulsed-field-ionization zero-kinetic-energy (PFI-ZEKE) photoelectron spectroscopy by recording transitions from the X̃ ^{3}B_{1} (0,0,0) ground vibronic state of CH_{2} which has been produced by photolysis of CH_{2}CO at 351 nm. The rotational energy level structure and the spin-rotation splittings of levels with K_{a} ≥ 1 could be resolved. The rotational structure in the X̃ ^{2}A_{1} (0, 0, 0) state deviates strongly from that of a rigid rotor. The adiabatic ionization potential of ^{3}CH_{2} has been determined to be (83772±3) cm^{−1}. Analysis of the photoionization selection rules indicates that the photoelectron is ejected as a superposition of ℓ even and ℓ odd partial waves.